# MOOSE-Chem2: Exploring LLM Limits in Fine-Grained Scientific Hypothesis Discovery via Hierarchical Search

### v1

@@ 3.3 Q2: Whether Hypotheses Judged Better by LLMs Exhibit Stronger Alignment With Ground-Truth Hypotheses?
Soft Recall Hard Recall
Greedy Search 16.60% 9.90%
w/ Self-consistency 31.50% 17.70%
HHS 40.40% 23.00%
Table 2: Recall of ground-truth components by discovered hypotheses.
@@
Given the absence of established metrics for this task, we propose an LLM-based evaluation that quantifies how well discovered hypotheses recall key chemical components of the ground-truth. Each hypothesis is decomposed into components regarding to the research question, and recall is assessed for each ground-truth component on a 0–3 scale, where 0 denotes no recall and 3 denotes an exact match.
We report two metrics: (1) Soft Recall, which counts a component as recalled if its score is greater than 0, normalized by the total number of components; (2) Hard Recall, which sums the raw 0–3 scores and normalizes them by the maximum possible score.
As shown in Table 2, the hypotheses discovered by HHS (which are preferred by LLMs, as demonstrated in § 3.2) achieve consistently higher recall with the ground-truth hypotheses.
3.4 Q3: Whether Defining the Reward Landscape With an Ensemble of Diverse LLMs of Similar Capacity Yields Better Outcomes Than Using Multiple Instances of the Strongest LLM Within That Group?
@@
Table 3: “EF”: Effectiveness, “NV”: Novelty, “DT”: Detailedness, “FS”: Feasibility, “OV”: Overall. “(GT)” and “(GM)” indicate that the pairwise evaluations were conducted by GPT-4o-mini and Gemini-1.5-flash, respectively.
The hypothesis optimization in HHS depends on the “$h_{\text{cur}}>h_{\text{prev}}?$” module (Figure 2), which provides the gradient signal. This raises the question: does a diverse ensemble of similarly capable LLMs enhance search performance compared to multiple instances of its strongest model?
To answer this, we design three experimental settings: (1) Mixed Committee: the “$h_{\text{cur}}>h_{\text{prev}}?$” module is implemented by an ensemble of three different LLMs—GPT-4o-mini (OpenAI, 2024), Gemini-1.5-flash (Georgiev et al., 2024), and Claude-3-haiku (Anthropic, 2024); (2) GPT-4o-mini Committee: the module is implemented by three instances of GPT-4o-mini; (3) Gemini-1.5-flash Committee: the module is implemented by three instances of Gemini-1.5-flash. All three settings use GPT-4o-mini as the proposer module for generating edits $d$ at each hierarchy level $i$.
@@
As shown in Table 3, across both evaluators, the GPT-4o-mini committee consistently outperforms the mixed committee, which in turn outperforms the Gemini-1.5-flash committee. These results suggest that leveraging repeated instances of the strongest single model provides a more effective gradient for hypothesis optimization than combining different models of similar capacity.
3.5 Q4: Do Multiple Identical LLMs Provide a Better Reward Landscape Than a Single LLM?
@@
Table 4: Pairwise comparison between HHS-1 and HHS-3.
In prior experiments (Tables 1 and 2), the reward landscape was defined using an ensemble of three identical LLMs, followed by a fourth instance of the same LLM that aggregated the three judgments into a final, reasoned decision. In the experiment corresponding to Table 3, the reward landscape was defined using an ensemble of three diverse LLMs of similar capacity, with aggregation again performed by a fixed instance of GPT-4o-mini. However, the extent to which this ensemble-based reward landscape improves performance compared to using a single LLM remains unclear.
To evaluate this, we compare two variants: HHS-3, which uses an ensemble of three identical instances of GPT-4o-mini to provide the reward signal, and HHS-1, which relies on a single instance of the same model.
Table 4 reports LLM-based pairwise evaluations between the two setups across four criteria. While overall quality, effectiveness, and detailedness are largely comparable, HHS-3 outperforms in novelty, whereas HHS-1 shows an advantage in feasibility.
This result is somewhat counterintuitive. Although both HHS-3 and HHS-1 use the same base LLM, they differ in how the gradient signal—that is, the decision of whether a new hypothesis improves upon the current one—is computed. In HHS-3, three independent comparative judgments are sampled using the same LLM, and a fourth instance of the same model aggregates these judgments by evaluating the underlying rationales and selecting the most justified preference.
Crucially, this summarization step is not a simple majority vote. Instead, the LLM is explicitly instructed to assess the relative strength of reasoning across all three perspectives and to favor the most compelling argument. This setup implicitly allows the model to surface and validate minority-supported but well-reasoned views, enabling exploration of more novel or unconventional hypotheses.
As a result, the aggregated gradient signal in HHS-3 may be more receptive to creative or atypical ideas that would otherwise be dismissed in a single-shot comparison. In contrast, HHS-1 relies on only a single comparative judgment per step, which reflects a narrower and more internally consistent heuristic perspective. This tends to produce more conservative refinements—favoring feasibility and coherence, but often at the cost of novelty.
The observed trade-off—greater novelty in HHS-3 and greater feasibility in HHS-1—thus stems not from classical ensemble averaging, but from the summarizing LLM’s ability to integrate diverse heuristic signals in a principled manner during hypothesis evaluation.
Soft Recall Hard Recall
HHS-3 40.40% 23.00%
HHS-1 32.40% 19.90%
Table 5: Recall of groundtruth components of groundtruth hypotheses between HHS-1 and HHS-3.
@@
These results suggest that while the performance gains from ensembling are relatively modest, using multiple identical LLMs to define the reward landscape may yield a more reliable and diverse evaluation signal than relying on a single instance.
4 Related Work
LLM-driven scientific discovery methods typically fall into two categories: (1) direct generation of hypotheses from a research background—comprising a research question and established methodologies (Qi et al., 2023); or (2) retrieval of seemingly unrelated but potentially useful knowledge fragments—termed inspirations—which are then combined with the background to construct a hypothesis (Yang et al., 2024a, b; Wang et al., 2024; Liu et al., 2025). While these methods have shown promise in generating novel ideas, they are often criticized for producing hypotheses that are overly coarse, lacking in detail, or omitting actionable experimental steps (Wang et al., 2024; Hu et al., 2024; Si et al., 2024). In contrast, our goal is to investigate how LLMs can be leveraged to generate fine-grained scientific hypotheses—those sufficiently detailed to be directly implemented in laboratory settings. Notably, inspiration-based hypotheses from prior work (Yang et al., 2024a, b) can serve as inputs to our framework in the form of coarse-grained hypothesis directions. These hypotheses emphasize novelty, as they often draw from previously unassociated knowledge, while our work complements them by enhancing experimental specificity and validity.
@@ 5 Conclusion
We present the first systematic study of fine-grained scientific hypothesis discovery with LLMs, framing it as a combinatorial optimization problem. To address the vast, implicit search space, we propose hierarchical heuristic search (HHS), which incrementally refines hypotheses from coarse to fine-grained levels. HHS smooths the reward landscape, stabilizes optimization, and consistently finds higher-quality hypotheses than flat search methods. Experiments show that (1) HHS reliably discovers better local optima than baselines, (2) LLM-preferred hypotheses align more closely with expert ground truths, and (3) repeated use of the strongest model provides better reward landscapes than diverse ensembles. These results underscore the value of hierarchical search in exploring more LLMs’ potential for scientific discovery.
Appendix A Expert Evaluation Instructions
@@
Appendix B Case Study
@@
Appendix C Hypothesis Search Prompt
The black prompt are shared prompt between the baseline methods and HHS, and the orange prompt is HHS specific prompt describing hierarchy. Below is the prompt:
@@ Appendix C Hypothesis Search Prompt
Appendix D Experiment Compute Resources

### v2

@@ 3.3 Q2: Whether Hypotheses Judged Better by LLMs Exhibit Stronger Alignment With Ground-Truth Hypotheses?
Table 2: Recall of ground-truth components by discovered hypotheses. #Steps represents the number of reasoning steps used. HHS represents HHS-3 referred in § 3.5.
Soft Recall Hard Recall #Steps
ChemCrow (M. Bran et al., 2024) 12.28% 7.20% -
Qi et al. (2024) 19.57% 11.15% -
SciMON (Wang et al., 2024) 18.57% 10.09% -
MOOSE (Yang et al., 2024) 20.04% 11.76% -
MOOSE-Chem (Yang et al., 2025) 19.99% 11.98% -
Greedy Search 16.60% 9.92% 9.69
   w/ In-context RL 16.76% 10.28% 16.80
   w/ Self-consistency 31.53% 17.73% 67.55
HHS (HHS-3) 40.35% 23.04% 282.04
   w/ In-context RL 33.63% 21.29% 531.08
   w/ Single LLM Gradient (HHS-1) 32.40% 19.95% 747.92
Given the lack of established metrics for this task, we introduce an LLM-based evaluation that measures how well the discovered hypotheses recover the methodological and experimental details of the ground-truth hypotheses. The detailed formulations of the two metrics—Soft Recall and Hard Recall—are provided in Appendix C.
As shown in Table 2, the hypotheses discovered by HHS—corresponding to better local optima than those produced by greedy search baselines—achieve consistently higher recall scores than both greedy search baselines and other comparative methods. Here, in-context RL refers to a mechanism in which, if the current hypothesis $h_{\text{cur}}$ does not outperform the previous one $h_{\text{prev}}$, $h_{\text{cur}}$ is inserted into the LLM’s context to generate a new candidate. We also report the total number of reasoning steps used by each method. A general trend emerges: increasing the number of reasoning steps tends to improve recall up to a point, beyond which excessive steps lead to diminishing or negative returns.
3.4 Q3: Whether Defining the Reward Landscape With an Ensemble of Diverse LLMs Yields Better Outcomes Than Using the Same Number of the Strongest LLMs Among Them?
Table 3: “EF”: Effectiveness, “NV”: Novelty, “DT”: Detailedness, “FS”: Feasibility, “OV”: Overall. “(GT)” and “(GM)” indicate that the pairwise comparisons were conducted by GPT-4o-mini and Gemini-1.5-flash, respectively.
@@
The optimization in HHS relies on the $h_{\text{cur}}>h_{\text{prev}}?$” module (Figure 2), which acts as the gradient signal driving the search. This raises a key question: does a diverse ensemble of comparably capable LLMs improve search performance, or is it more effective to use the same number of parallel instances of the single strongest LLM among the ensemble?
To answer this, we design three experimental settings: (1) Mixed Committee: the $h_{\text{cur}}>h_{\text{prev}}?$” module is implemented by an ensemble of three different LLMs—GPT-4o-mini (OpenAI, 2024), Gemini-1.5-flash (Georgiev et al., 2024), and Claude-3-haiku (Anthropic, 2024); (2) GPT-4o-mini Committee: the module is implemented by three instances of GPT-4o-mini; (3) Gemini-1.5-flash Committee: the module is implemented by three instances of Gemini-1.5-flash. Each committee’s three judgments are then aggregated by a GPT-4o-mini, which produces the final decision for $h_{\text{cur}}>h_{\text{prev}}?$” representing that committee. All three settings use GPT-4o-mini as the proposer module for generating edits $d$ at each hierarchy level $i$.
@@
As shown in Table 3, across both evaluators, the GPT-4o-mini committee consistently outperforms the mixed committee, which in turn outperforms the Gemini-1.5-flash committee. These results suggest that leveraging repeated instances of the strongest single model provides a more effective gradient signal for hypothesis optimization than combining different models of similar capacity.
3.5 Q4: Do Multiple Identical LLMs Yield a Better Reward Landscape Than One?
In experiments of Tables 1 and 2 (except for the HHS-1 line in table 2), the reward landscape was defined using an ensemble of three identical LLMs, followed by a fourth instance of the same LLM that aggregated the three judgments into a final, reasoned decision. However, it is unclear on whether one LLM would be already enough. To evaluate this, we compare two variants: HHS-3, which uses an ensemble of three identical instances of GPT-4o-mini (and a fourth instance of the same LLM for aggregation) to provide the reward signal, and HHS-1, which relies on a single instance of the same model. Table 4 reports LLM-based pairwise evaluations between the two setups across four criteria. While overall quality, effectiveness, and detailedness are largely comparable, HHS-3 outperforms in novelty, whereas HHS-1 shows an advantage in feasibility.
Table 4: Pairwise comparison between HHS-1 and HHS-3 with LLM-based evaluation.
@@ 3.5 Q4: Do Multiple Identical LLMs Yield a Better Reward Landscape Than One?
This result is somewhat counterintuitive. Notably, the summarization step is not a simple majority vote: the aggregating LLM assesses the relative strength of reasoning across the three perspectives and selects the most compelling argument. This allows it to surface minority-supported but well-reasoned views, promoting exploration of more novel or unconventional hypotheses. Consequently, the aggregated reward signal in HHS-3 becomes more sensitive to creative or atypical ideas that a single-shot evaluation might overlook, whereas HHS-1 relies on a single comparative judgment at each step—favoring conventional and well-established hypotheses, at the expense of novelty.
@@
4 Related Work and Discussion
LLM-driven scientific discovery methods typically fall into two categories: (1) direct generation of hypotheses from a research background—comprising a research question and background survey (Qi et al., 2024); or (2) retrieval of seemingly unrelated yet potentially useful knowledge fragments, or inspirations, which are then combined with the background to construct a hypothesis (Yang et al., 2024, 2025; Wang et al., 2024; Liu et al., 2025b). While these methods show promise in generating novel ideas, they are often criticized for producing hypotheses that are overly coarse, lacking detail, or omitting actionable experimental steps (Wang et al., 2024; Hu et al., 2024; Si et al., 2025). In contrast, our goal is to investigate how LLMs can generate fine-grained scientific hypotheses—those sufficiently detailed to be directly implemented in laboratory settings. Although this work centers on the pre-experimental stage of discovery, the framework can in principle extend to the experiment-guided stage (Liu et al., 2025a; Romera-Paredes et al., 2024; Shojaee et al., 2025; Novikov et al., 2025) by incorporating experimental feedback into the background survey. Likewise, details retrieved from papers relevant to the proposed hypothesis can also be integrated into the background survey.
@@ 5 Conclusion
We introduce and formalize the fine-grained scientific hypothesis discovery task as a combinatorial optimization problem. To explore the upper limit of LLMs’ capacity for this task, we frame it as an optimization problem, and propose hierarchical heuristic search (HHS), which theoretically smooths the reward landscape, reduces combinatorial complexity through optimal substructure exploitation, and identifies superior local optimum by interpolating among discovered optima. Empirical results show that (1) HHS reliably discovers better local optima than flat greedy search baselines, (2) hypotheses preferred by LLMs often align more closely with expert annotations, (3) repeated use of the strongest model defines a more effective reward landscape than diverse ensembles, and (4) aggregating identical LLMs yields a reward signal more sensitive to novelty and higher in recall than single-instance evaluation. These findings illustrate how hierarchical search can better harness LLMs’ internal heuristics for scientific discovery. Although evaluated on a chemistry dataset, the proposed task formulation and methodology are expected to generalize across disciplines, with only the manually designed hierarchy (Figure 1) requiring domain-specific adaptation.
Acknowledgments
This work was supported by a locally commissioned task from the Shanghai Municipal Government. This work is supported by Shanghai Artificial Intelligence Laboratory. This research/project is supported by the Ministry of Education, Singapore under its MOE Academic Research Fund Tier 2 (STEM RIE2025 Award MOE-T2EP20123-0005).
Appendix A MOOSE-Chem2 Overview I/O Figure
Figure 5: Overview of the input and output of the Hierarchical Heuristic Search (HHS) framework, also known as MOOSE-Chem2.
Figure 5 presents an overview of the input and output of the Hierarchical Heuristic Search (HHS) framework, also referred to as MOOSE-Chem2. The input, a coarse-grained research direction, may range from a brief sentence outlining a general research direction to a hypothesis that already includes many methodological and experimental details—such as one produced by the MOOSE-Chem (Yang et al., 2025) framework.
Appendix B Expert Evaluation Instructions
@@
Appendix C Recall Metric Calculation
Each evaluation compares a pair of hypotheses—a ground-truth hypothesis (from literature) and a candidate hypothesis (generated by the model). The goal is to quantify how much methodological or experimental content from the ground-truth hypothesis is covered by the candidate.
Step 1. Decomposition.
Both hypotheses are decomposed into methodological and experimental components (details) using structured prompts co-designed with chemistry PhD students. The prompts guide an LLM to extract key components such as procedures, specific materials, or reaction parameters. Each component is annotated with its role, function, and context.
Step 2. Scoring via LLM-Judge.
For each ground-truth component, the LLM judge searches for functionally corresponding components in the candidate hypothesis and assigns a coverage score from 0–3:
  • 0: no match;
  • 1: vague or partial match;
  • 2: close but not exact match;
  • 3: exact or specific match.
The scoring prompts and examples were jointly developed and audited by PhD students to ensure calibration and consistency.
Step 3. Aggregation.
Two recall metrics are computed:
Soft Recall $=\frac{\#\{\text{components with score}>0\}}{\#\{\text{ground-truth components}\}},$ (7)
Hard Recall $=\frac{\sum_{i}\text{score}_{i}}{3\times\#\{\text{ground-truth components}\}}.$ (8)
Soft Recall captures whether relevant components are recovered at all, while Hard Recall further reflects how precisely they are reproduced.
Appendix D Case Study
@@
Appendix E Expert Analysis of Hypothesis Quality
E.1 Convergence to Ground-Truth Local Optima
To complement our quantitative evaluations, we asked domain experts to qualitatively assess how well the hypotheses generated by HHS aligned with the expert-annotated fine-grained hypotheses in our benchmark.
The distribution of expert assessments across all evaluated examples is as follows:
  • Reached a completely different region—likely a distinct local optimum with scientific plausibility: 60%
  • Reached the vicinity of the ground-truth local optimum, but with differing details: 24%
  • Reached the vicinity of the ground-truth local optimum, but failed to fully elaborate or specify the key details: 16%
Here, the “ground-truth local optimum” refers to the expert-extracted fine-grained hypothesis from a publication, which serves as the reference target. “Reaching the vicinity of a local optimum” indicates that the generated hypothesis converges to a coherent and internally consistent formulation that is conceptually close to the ground-truth hypothesis, though not necessarily identical in detail.
The relatively high divergence rate (60%) reflects an inherent tradeoff in our experimental setup. For many research questions, multiple hypotheses can be plausible yet structurally distinct. Guiding the model toward the exact ground-truth hypothesis requires:
  1. 1.
    initializing the search process from a starting point sufficiently close to the ground-truth optimum;
  2. 2.
    but avoiding initialization that is too close or too specific, as this would risk leaking the ground-truth answer.
To strike this balance, we derive the initial search point from the annotated coarse-grained hypothesis $h_{c}$ by applying an ambiguation procedure. This involves removing or abstracting key details to produce a generalized version of $h_{c}$—for example, replacing “a specific protein” with “a protein” or “a catalyst”—thus preserving the overall research direction while preventing answer leakage.
Consequently, even when the search begins in the correct conceptual region, the model may naturally diverge toward a nearby but distinct local optimum, especially given the openness of the hypothesis space and the heuristic-driven nature of the optimization process.
E.2 Coverage of Experimentally Critical Details
In addition to alignment with the reference hypotheses, we evaluated the extent to which the generated hypotheses captured the critical experimental details required for practical implementation.
Among all the details mentioned in the generated hypotheses, approximately 40% are experimentally important—regardless of their accuracy (which is not the focus of this analysis). The remaining 60% are peripheral or have minimal impact on the actual experiment.
Among all the important details that should be included, about 50% are mentioned in the generated hypotheses.
Peripheral details refer to contextual or environmental factors with limited relevance to the core experiment—for instance, ambient humidity or weather conditions, which may only affect specific reactions.
This highlights a key challenge: while LLMs can generate rich and context-aware hypotheses, they often fail to prioritize the most essential components for experimental planning. Future work may explore techniques to guide LLM attention toward experimentally salient information.
Table 5: Error analysis of hypotheses generated by HHS. Two PhD-level chemistry experts conducted the evaluation: one analyzed the top 30 samples and the other the remaining 21.
Missing key chemical substances 14/30
Excessive details in characterization methods 28/30
Feasibility issues 18/30
Limitations of characterization methods 08/30
Insufficient basis for material selection 22/30
Lack of design comparison experiments 12/30
Ignoring data validation and reproducibility 10/30
Severe deviation from feasibility 8/21
Missing or incorrect key chemicals or reaction systems 9/21
Incorrect explanation of chemical principles 12/21
Incorrect prediction of experimental system 10/21
Table 6: Error analysis on why HHS’s hypotheses are better than the greedy search baselines’. Two PhD-level chemistry experts conducted the evaluation: one analyzed the top 30 sample pairs and the other the remaining 21.
Insufficient performance metrics 25/30
Complexity of experimental conditions 16/30
Insufficient explanation details 29/30
Inadequate description of preparation plan 22/30
Vague research objectives 28/30
Cost and scalability issues 13/30
Poor feasibility 12/21
Errors in research plan details 21/21
Insufficient explanation details 19/21
Clear experimental system 21/21
Appendix F Error Analysis
Two PhD-level chemistry experts conducted the error analyses. Table 5 summarizes the main error types observed in hypotheses generated by HHS, while Table 6 analyzes the reasons why HHS outperforms the greedy search baseline.
Appendix G Hypothesis Search Prompt
The following prompt is used to guide both the baseline methods and our proposed method, HHS, during the hypothesis refinement process. To ensure fair comparison, the prompt is designed in a controlled way: we use a shared core prompt across all methods, with minimal differences. Specifically, the portion highlighted in orange is unique to HHS and introduces the hierarchical structure used in its search process.
This design isolates the effect of hierarchical search. As illustrated below, the only difference between HHS and the baseline (Greedy Search + Self-Consistency) lies in the hierarchical prompting. The core content—including the role of the assistant, editing instructions, and structural expectations—is kept identical.
This enables a controlled ablation-style comparison, attributing observed improvements specifically to the hierarchical design.
The complete prompt is as follows:
@@ Appendix G Hypothesis Search Prompt
Appendix H Experiment Compute Resources